Graph the supply and demand curves. Be sure to calculate the P and Q intercepts for demand and the P intercept for supply. Calculate and illustrate the equilibrium price and quantity. [Hint: Show your work. [5]
1
Expert's answer
2021-10-13T05:36:02-0400


Assuming we have quantity demanded and supplied at the given prices on the table below which is a representation of the rice f gasoline.






A representation of the data on a graph, produces the the demand and supply curve that appears as shown below.







Calculations

Step 1


Determining the demand slope (Since it is a curve we have to establish the gradient)

in this case the gradient points are (500, 1.60) and (800, 1.00)


"=" "\frac{dQ_2-dQ_1}{dP_2-dP_1}"


"=" "\frac{500-800}{1.60-1.00}"


"=" "\frac{-300}{0.60}" "=" -500


"b = -500"


Demand equation


"Q_d = a + bP"


Where:

b"-" is the slope

"-" is quantity demanded when price is zero


replacing the equation with numbers from the table above to determine "a"


"800 = a - 500 (1.00)"

"a = 800 + 500"

"a = 1,300"


therefore the demand equation is


"Qd = 1,300 - 500P"


Step 2

Determining the supply slope (Since it is a curve we determine the gradient)

In this case the points of the gradient are (800,2.20) and (500, 1.00)


"=" "\frac{dQ_2 - dQ_1}{P_2 - P_1}"


"=" "\frac{800-500}{2.20-1.00}"


"=" "\frac{300}{1.20}"


"b = 250"


Supply equation


"Q_s = a + bP"


Where:

b"-" is the slope

"-" is quantity supplied when price is zero


replacing the equation with numbers from the table above to determine "a"


"500 = a + 250 (1.00)"


"a = 500 - 250"


"a = -250"

therefore the supply equation is


"Q_s = -250 + 250P"


At equilibrium Supply equals demand


"Q_s = Qd"


"-250 + 250P = 1,300 -500P"


"500P + 250P = 1,300 - 250"


"750P = 1,050"


Therefore the equilibrium price is


"P_e = 1.40"


Determining the equilibrium quantity


"Q_s = Q_d = Q_e = 1,300 - 500 (1.40)"


"Q_e = 1,300 - 700"


"Q_e = 600"